OpenStudy (monica0597):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (ghazi):

(y-4)= (0-4)/(2+6)* (x+6)

OpenStudy (ghazi):

\[(y-4)=\frac{ (0-4) }{ 6+2 }*(x+6)\]

OpenStudy (ghazi):

use two point slope form equation

OpenStudy (monica0597):

what i get is y − 4 = 1/2(x + 6) and it isnt one of the choices...

OpenStudy (ghazi):

yes that's correct but slope is -1/2

OpenStudy (monica0597):

ohhh... thanks :)

OpenStudy (ghazi):

:)